# Rethinking Prompt Optimizers: From Prompt Merits to Optimization

### v1

@@ 3.2 Training the Merit-Guided Prompt Optimizer
To guide the model toward generating more preferred prompts, we train the optimizer using the Direct Preference Optimization (DPO) framework (Rafailov et al., 2023). The objective is to learn a prompt optimizer that prefers merit-aligned prompts yielding higher-quality responses.
@@
where $M_{o}$ denotes the optimization model, and $M_{\text{ref}}$ is a reference model with adapters disabled.
@@ 4.1 Setups
MePO is trained using Qwen2.5-7B-Instruct4, selected for its strong generation ability among similarly sized models Qwen et al. (2025). This choice ensures that the optimizer is built on a lightweight LLM that maximizes performance within its size class—allowing efficient deployment without sacrificing generation quality.
Datasets We evaluate our method on five benchmark datasets: ARC-Easy, ARC-Challenge (Clark et al., 2018), GSM8K (Cobbe et al., 2021), BigBench-Hard (BBH) (Suzgun et al., 2023), and PiQA (Bisk et al., 2020). Following Lu et al. (2025), we adopt 3-shot for GSM8K, BBH, and PiQA, and zero-shot for the ARC. Prompt optimization is applied only to the test query, while the in-context examples remain unchanged.
@@ 4.2 Main Results
Efficacy MePO consistently outperforms all baselines across datasets and task models, with average accuracy gains of 1.48% on Qwen2-7B, 2.40% on Tulu2-7B, 2.58% on LLaMA2-7B, and 3.92% on LLaMA3-8B. Notable improvements are observed on ARC-Challenge with LLaMA3-8B (+6.91%), GSM8K with LLaMA2-7B (+5.4%), and ARC-Easy and BBH with Tulu2-7B (+4.5%, +3.83%).
@@
Qwen2-7b Training-Free 82.79 68.17 81.52 49.65 82.32 72.89
MePO 83.33 68.52 83.12 53.42 83.46 74.37
Tulu2-7b-dpo Training-Free 49.58 34.30 32.93 37.66 72.60 45.41
MePO 55.05 38.14 35.18 39.89 73.61 48.37
LLaMA2-7b-chat-hf Training-Free 36.91 30.55 20.91 36.60 53.10 35.61
MePO 39.86 31.74 25.25 38.68 55.33 38.17
LLaMA3-8b-instruct Training-Free 52.36 50.77 58.25 39.58 71.69 54.53
MePO 55.01 51.88 59.59 42.19 76.66 57.07
@@ 4.3 Ablation Study: Training-Free Comparison
As summarized in Table 1, our filtering process retains 25.5k samples from the original 30k Alpaca entries and 10k from 14k BPO entries, discarding cases where merit-optimized prompts failed to outperform their raw counterparts. This ensures that MePO is trained only on successful optimization examples, encouraging the model to learn effective and reliable merit application.
To determine whether MePO’s performance gains result purely from the use of high-quality prompt merits or from learning how to apply them effectively, we compare it against a training-free baseline. This baseline uses our identified merits as static templates, without any learned optimization. As shown in Table 3, MePO consistently outperforms the training-free variant across all datasets and model backbones. Notably, average accuracy improves by 2.96% on Tulu2 and over 2.5% on LLaMA2 and 3, demonstrating the value of end-to-end optimization over training-free prompting.
@@ 4.4 Case Study: Upward Compatibility
LLaMA2-13b-chat-hf Tulu2-70b-dpo
- MePO - MePO
ARC-Easy 49.83 55.93 80.47 85.23
ARC-Challenge 45.05 52.73 58.87 60.49
GSM8K 28.73 35.97 61.41 64.67
BBH 36.48 43.72 48.05 49.74
PiQA 56.13 60.99 77.75 80.14
@@
As shown in Table 4, applying MePO prompts to larger inference LLMs consistently improves performance across all datasets. These results indicate that MePO—despite being a lightweight prompt optimizer—generates clear and concise reasoning guidance that are well interpreted by large-scale LLMs, demonstrating strong upward compatibility.
@@ 5 Related Work
Automatic Prompt Optimization (APO) offers a practical alternative to fine-tuning by focusing on prompt engineering instead of parameter updates. Beyond exploration-based methods search over predefined prompt pools (Ma et al., 2023; Shi et al., ), recent works favor discrete exploitation-based methods that directly generate or revise prompts without relying on the quality of initial pools, making them more flexible and better suited for generative tasks beyond fixed-answer classification.
Online-Based Discrete APO Several works perform discrete prompt optimization using high-capacity, API-based LLMs as optimization models. For example, Zhou et al. (2022) use InstructGPT; Wang et al. ; Guo et al. (2024); Pryzant et al. (2023) rely on GPT-3.5; Xiang et al. (2025) adopt Claude 3.5 Sonnet; and Yuksekgonul et al. (2024) utilize GPT-4o. While effective, these approaches depend on proprietary APIs, introducing practical concerns around cost, inference latency, and privacy risks.
@@
Prompt Merits Exploration To improve prompt quality, Arora et al. shows that prompts in open-ended QA formats outperform restrictive ones. Wei et al. (2022) introduces intermediate reasoning steps in prompts. Zhou et al. (2023) propose decomposing prompts into simpler subcomponents. Besides, several studies provide design principles for prompt optimization (Bsharat et al., 2023; Ye et al., 2024). However, these merits are largely proposed heuristically and validated through downstream experiments, leading to inconsistencies in perspective. In contrast, we derive prompt merits by prompt–response empirical analysis, grounding them in measurable improvements.
@@ Appendix A Empirical Study: Prompt Optimization Across Model Scales
Evaluating GPT-4 Optimized Prompts on Smaller Inference Models.
@@
These examples show that prompts crafted by powerful optimization models may not be directly interpretable by smaller models due to verbosity or overly rich CoT structure, leading to inconsistent outputs.
Evaluating Lightweight LLMs with Complex Optimization Algorithms.
We assess whether lightweight LLMs can follow optimization prompts designed for advanced models. Specifically, we use the Differential Evolution (DE) algorithm from Guo et al. (2024), which guides prompt generation using EvoPrompt instructions. We supply the EvoPrompt Prompt and DE-generated Optimal Prompt as input to the model and use DeepSeek-R1 to compare their quality, as shown in Fig. 5.
@@
We further test whether lightweight LLMs can act as effective optimization models when guided by clear, interpretable merits. Using the Basic Prompt from Guo et al. (2024), we apply our proposed merit-based instruction (see Fig. 14) to generate optimized prompts. We then evaluate the difference between Basic and Optimized Prompts using DeepSeek-R1 with the pairwise setup shown in Fig. 9.
@@ Appendix B Identifying Effective Prompt Merits via Comparative Evaluation
To explore the characteristics of effective prompts, we conduct a comparative evaluation using a series of specially designed prompts.
@@
We begin by using a rewriting prompt (shown in Fig. 6) to generate multiple reformulations of each raw question. Specifically, five rewritten versions are produced for each input prompt to preserve diversity. Below is an example from the Alpaca dataset, where the raw prompt results in an incorrect response, while Rewrite 1 and Rewrite 2 yield correct answers under the same inference model4:
@@ Appendix C POP Dataset Construction and Training Details
To simulate noisy user inputs—such as “how can i go fr sigapor?”—we degrade 10% of the raw prompts from the Alpaca and BPO datasets using a base model. The prompt used for degradation is shown in Fig. 7. Below is an example from the Alpaca dataset, generated by the model referenced in 4:
@@
We fine-tune MePO based on Qwen2.5-7B-Instruct4 using 4 NVIDIA RTX A5000 GPUs. Each input instruction $x$ is truncated or padded to 2000 tokens. Training hyperparameters and cost are summarized in Table 6.
@@
Appendix D Dataset Information and Experimental Implementation
Evaluation Protocol.
GSM8K, PiQA, and BBH are evaluated in a 3-shot setting to facilitate answer extraction, while all other datasets are evaluated in a zero-shot setting. For consistency, all BBH tasks are reformatted into multiple-choice format, except for multistep_arithmetic_two, object_counting, and word_sorting, which retain their original formats. PiQA, which includes two-solution questions, is also converted to multiple-choice format. For all multiple-choice datasets, we determine the answer by selecting the option (e.g., A, B or 1, 2) with the highest next-token logit probability following the ##Answer: token.
Prompt Templates.
Fig. 12 shows the prompt formats used for GSM8K, PiQA and ARC. For BBH, the multiple-choice tasks use the PiQA prompt while the math tasks use the GSM8K prompt.
Baseline Implementations.
@@ Appendix D Dataset Information and Experimental Implementation
LLaMA3.2-3b-instruct - 55.46 73.34 63.23 35.6 66.32 58.79
Inference Model 60.75 78.16 63.00 37.91 69.37 61.84
BPO 55.74 74.75 56.71 37.34 67.14 58.34
FIPO 67.66 78.66 65.13 37.65 69.53 63.73
MePO 61.43 79.58 65.66 40.49 70.29 63.49
@@ D.1 Case Study: Downward Compatibility
As shown in Table 7, both BPO and FIPO exhibit performance degradation when paired with lightweight inference models. For example, under LLaMA-3.2-3B-Instruct, BPO drops by 5.01% on ARC-Easy, 3.41% on ARC-Challenge, and 2.23% on PiQA. Under LLaMA-3.2-1B-Instruct, FIPO drops by 3.60% on ARC-Easy, and BPO drops by 4.39% on GSM8K. Both also show performance decreases on BBH under both lightweight inference models. In contrast, MePO consistently improves performance across all datasets, demonstrating strong downward compatibility and robustness in low-resource settings.
@@ D.2 Case Study: Online-Based Discrete APO Comparison
To further validate MePO’s effectiveness in optimizing prompts or instructions with refined meta-prompts, we compare MePO with three discrete APO methods:
@@
Table 8: Meta-prompts used for each baseline discrete APO method.
@@ D.3 BBH Experimental Results
date_understanding 44.8 30.8 28.8 31.6 42 25.6 27.2 62
disambiguation_qa 63.2 32.8 31.2 32.4 30 36.4 33.6 68.8
hyperbaton 75.6 48.4 48.4 48.4 48.4 51.6 48.4 78
logical_deduction_five_objects 49.2 22.4 18 28 33.2 24.4 19.6 50
logical_deduction_seven_objects 42.4 17.6 14.8 30 27.2 30.4 15.2 42.8
logical_deduction_three_objects 66.4 32.8 31.6 33.6 36.4 35.2 31.6 71.2
movie_recommendation 48 26.8 22.4 33.2 23.6 23.2 22.4 52.8
penguins_in_a_table 54.11 21.23 22.6 31.51 46.58 26.71 23.97 50.68
reasoning_about_colored_objects 47.6 32.4 32 16.4 29.2 30 16.4 46
ruin_names 37.2 24 30.8 35.2 27.2 41.6 28.4 66.8
salient_translation_error_detection 32 26.8 23.2 21.6 28 34 16.8 52
snarks 60.11 44.94 46.07 46.07 50.56 57.3 46.07 81.46
word_sorting 6 3.2 2.8 3.2 4.8 3.2 1.2 10.4
temporal_sequences 61.2 19.2 28.4 28.4 33.2 28 28.4 62
tracking_shuffled_objects_five_objects 28.4 19.2 20 20.8 30.8 20 20 17.2
tracking_shuffled_objects_seven_objects 13.2 24 14.4 13.6 22 11.2 12.8 13.6
tracking_shuffled_objects_three_objects 27.2 29.2 24.4 26.8 49.2 30 30.8 28
causal_judgement 62.57 49.2 51.87 51.34 49.2 47.59 51.87 53.48
formal_fallacies 54 52.8 53.2 53.6 52.4 51.6 53.2 52
navigate 57.6 53.6 42 41.6 41.6 41.2 42 52.8
web_of_lies 46.4 48.8 48.8 49.2 54 44.4 46.4 46.8
sports_understanding 70 50 46.4 70.4 54 53.6 43.2 45.6
boolean_expressions 34.8 56.8 52.4 54.8 51.2 48 54 30.4
multistep_arithmetic_two 33.6 12.4 9.6 36 14 46 16 16.8
object_counting 36 33.2 32.8 50.4 33.2 48.8 46.8 49.6
@@
date_understanding 57.77 36.4 35.6 49.2 52.4 34 31.2 50
disambiguation_qa 66.4 38.4 37.2 48.8 54.8 44.8 40.8 68.8
hyperbaton 81.6 54.4 50.4 50.4 54.4 56.8 48.4 62
logical_deduction_five_objects 58.8 27.2 24.4 47.2 42.4 30 23.6 52
logical_deduction_seven_objects 49.2 29.6 20.4 39.6 36.8 21.2 18.4 51.6
logical_deduction_three_objects 82 36.8 51.2 42.8 39.2 40.4 34.8 65.6
movie_recommendation 63.2 38.4 26.8 38.4 29.2 32 27.2 34
penguins_in_a_table 56.85 28.77 25.34 38.36 54.79 31.51 25.34 51.37
reasoning_about_colored_objects 56.8 35.6 38 23.6 41.6 26 18.4 58.8
ruin_names 38 27.6 41.6 42 28.8 45.6 28.4 73.6
salient_translation_error_detection 43.2 36.8 38.8 28.8 31.2 39.2 23.6 54.8
snarks 67.42 53.37 51.12 49.44 66.85 60.11 46.07 83.15
word_sorting 16.8 8.4 8 8 10.7 7.3 3.2 16.8
temporal_sequences 38 37.6 39.2 33.2 42 29.2 28.8 66.8
tracking_shuffled_objects_five_objects 31.2 39.6 33.2 35.2 32.8 20.4 23.2 21.2
tracking_shuffled_objects_seven_objects 15.2 30 26.4 16.4 28.4 18.8 15.2 18
tracking_shuffled_objects_three_objects 33.6 32.4 33.2 33.2 64.8 31.6 31.6 29.6
causal_judgement 49.2 56.68 51.87 51.79 56.15 62.57 52.41 59.36
formal_fallacies 51.2 57.2 54.8 53.2 52.8 53.2 53.2 59.2
navigate 61.6 58.8 44.4 44.8 48.4 47.2 42 60
web_of_lies 51.2 55.2 51.6 51.6 59.2 51.2 49.6 49.6
sports_understanding 77.6 67.2 54.8 77.2 61.2 58 46 49.2
boolean_expressions 48.4 61.2 54 54.4 54 52.8 55.6 36.4
multistep_arithmetic_two 45.6 14.4 14.4 41.2 16 57.6 18.4 18
object_counting 39.6 35.2 60.4 55.2 34 60.8 49.6 53.6

### v2

@@ 3.2 Training the Merit-Guided Prompt Optimizer
To guide the model toward generating more preferred prompts, we train the optimizer using the Direct Preference Optimization (DPO) framework (Rafailov et al., 2023). The objective is to learn a prompt optimizer that prefers merit-guided prompts yielding higher-quality responses.
@@
[…]
@@ 4.1 Setups
MePO is trained using Qwen2.5-7B-Instruct2, selected for its strong generation ability among similarly sized models Qwen et al. (2025). This choice ensures that the optimizer is built on a lightweight LLM that maximizes performance within its size class—allowing efficient deployment without sacrificing generation quality.
Datasets We evaluate our method on five benchmark datasets: ARC-Easy, ARC-Challenge (Clark et al., 2018), GSM8K (Cobbe et al., 2021), BigBench-Hard (BBH) (Suzgun et al., 2023), and PiQA (Bisk et al., 2020). Following Lu et al. (2025), we adopt 3-shot for GSM8K, BBH, and PiQA, and zero-shot for ARC, with instructions for answer extraction. Prompt optimization is applied only to test query, while others remain unchanged.
@@ 4.2 Main Results
Efficacy MePO consistently outperforms all baselines across datasets and task models, with average accuracy gains of 1% on Qwen2-7B, 2.37% on Tulu2-7B, 2.59% on LLaMA2-7B, and 3.91% on LLaMA3-8B. Notable improvements are observed on ARC-Challenge with LLaMA3-8B (+6.91%), GSM8K with LLaMA2-7B (+5.4%), and ARC-Easy and BBH with Tulu2-7B (+4.5%, +3.69%).
@@
Task Model Optimizer ARC-Easy ARC-Challenge GSM8K BBH PiQA Avg.
Qwen2-7b Training Free 82.79 68.17 81.52 52.93 82.32 73.55
PattoPO 83.33 68.52 83.12 54.35 83.46 74.56
Tulu2-7b-dpo Training Free 49.58 34.30 32.93 40.96 72.60 46.22
PattoPO 55.05 38.14 35.18 43.25 73.60 49.04
LLaMA2-7b-chat-hf Training Free 36.91 30.55 20.91 39.88 53.10 36.27
PattoPO 39.86 31.74 25.25 41.97 55.33 38.83
LLaMA3-8b-instruct Training Free 52.36 50.77 58.25 42.80 71.69 55.17
PattoPO 55.01 51.88 59.59 45.52 76.66 57.73
@@ 4.3 Ablation Study: Training-Free Comparison
As shown in Table 1, we retain 25.5k Alpaca and 10k BPO samples by discarding cases where merit-guided prompts fail to outperform their raw counterparts. This ensures MePO is trained only on reliably successful examples, promoting robust learning of merit application.
To isolate the effect of learning, we compare MePO with a training-free baseline where the inference model itself rewrites raw prompts using static merit-guided templates (Fig. 14), and then generates responses from the rewritten prompts. This comparison tests whether learning to apply merits is more effective than applying them heuristically.
As shown in Table 3, MePO consistently outperforms the training-free approach across all datasets and model backbones. Notably, average accuracy improves by 2.38% on Tulu2 and over 2.59% on LLaMA2 and 3, demonstrating the value of end-to-end optimization over training-free prompting.
@@ 4.4 Case Study: Upward Compatibility
LLaMA2-13b
-chat-hf
Llama-3.3-70B
-Instruct
Tulu2-70b
-dpo
- MePO - MePO - MePO
ARC-Easy 49.83 55.93 95.03 92.34 80.47 85.23
ARC-Challenge 45.05 52.73 92.24 91.04 58.87 60.49
GSM8K 28.73 35.97 91.74 92.49 61.41 64.67
BBH 39.49 46.74 66.25 68.25 51.63 53.07
PiQA 56.13 60.99 87.87 89.93 77.75 80.14
@@
As shown in Table 4, applying MePO prompts to larger inference LLMs consistently matches or improves performance across all datasets. These results demonstrate that MePO, despite being lightweight, produces clear and concise reasoning guidance that large-scale LLMs can effectively interpret, highlighting its strong upward compatibility and potential for general applicability.
@@ 5 Related Work
Automatic Prompt Optimization (APO) provides a practical alternative to fine-tuning by optimizing prompts (Yang et al., ; Zhu et al., 2023). Beyond exploration-based methods search over predefined prompt pools (Ma et al., 2023; Shi et al., ; Zhu et al., 2025), recent works favor discrete exploitation-based methods that directly generate or revise prompts without relying on the quality of initial pools, making them more flexible and better suited for generative tasks beyond fixed-answer classification.
Online-Based Discrete APO Several works perform discrete prompt optimization using high-capacity, API-based LLMs as optimization models. For example, Zhou et al. (2022) use InstructGPT; Wang et al. (2024); Guo et al. (2024); Pryzant et al. (2023) rely on GPT-3.5; Xiang et al. (2025) adopt Claude 3.5 Sonnet; and Yuksekgonul et al. (2024) utilize GPT-4o. While effective, these approaches depend on proprietary APIs, introducing practical concerns around cost, inference latency, and privacy risks.
@@
Prompt Merits Exploration To improve prompt quality, Arora et al. show that prompts in open-ended QA formats outperform restrictive ones. Wei et al. (2022) introduce intermediate reasoning steps in prompts. Zhou et al. (2023) propose decomposing prompts into simpler subcomponents. Besides, several studies provide design principles for prompt optimization (Bsharat et al., 2023; Ye et al., 2024). However, these merits are largely proposed heuristically and validated through downstream experiments, leading to inconsistencies in perspective. In contrast, we derive prompt merits by prompt–response empirical analysis, grounding them in measurable improvements.
@@ Appendix A Empirical Study: Prompt Optimization Across Model Scales
A.1 Evaluating GPT-4 Optimized Prompts on Smaller Inference Models
To evaluate the downward compatibility of large-scale optimization models with lightweight inference models, beyond the results shown in Fig. 1(a), we conduct additional experiments that yield irrelevant responses.
@@ A.1 Evaluating GPT-4 Optimized Prompts on Smaller Inference Models
DeepSeek-R1 Evaluation: The response does not include irrelevant content, but the additional explanatory text is somewhat unnecessary for the task.
@@
DeepSeek-R1 Evaluation: High accuracy in the haiku itself but a minor deduction for the slightly off-topic preface.
These examples show that prompts crafted by powerful optimization models may not be directly interpretable by smaller models due to verbosity or overly rich CoT structure, leading to unnecessary or off-topic outputs.
A.2 Evaluating Lightweight LLMs with Complex Optimization Algorithms
We evaluate whether lightweight LLMs can follow optimization instructions designed for advanced, large-scale models. Specifically, we adopt the Differential Evolution (DE) algorithm from Guo et al. (2024), which uses EvoPrompt instructions to guide prompt generation. We provide both the EvoPrompt prompt and the DE-generated optimal prompt as input to lightweight LLMs, and compare their quality using DeepSeek-R1 (see Fig. 5).
As shown in Fig. 1(b), DeepSeek-R1 evaluations report that the EvoPrompt prompt is “clearer and more structured,” directly aligning with the task, while the prompt generated by the lightweight LLM “introduces unnecessary steps that may distract from the core goal.” These results indicate that GPT-3.5, as used in EvoPrompt, produces higher-quality prompts than those generated by LLaMA-3.2-3B or Tulu2-7b-dpo under the same algorithm—suggesting that lightweight LLMs struggle to serve as effective optimizers when guided by complex optimization instructions.
@@ A.2 Evaluating Lightweight LLMs with Complex Optimization Algorithms
We further test whether lightweight LLMs can act as effective optimization models when guided by clear, interpretable merits. Using the Basic Prompt from Guo et al. (2024), we apply our proposed merit-based instruction (see Fig. 14) to generate optimized prompts. We then evaluate the difference between Basic and Optimized Prompts using DeepSeek-R1 with the pairwise setup shown in Fig. 9. The same evaluation is conducted with GPT-4o under the same setup, and it produces consistent results, confirming that the merit-guided prompts generated by lightweight LLMs outperform the Basic Prompt in quality.
@@ Appendix B Identifying Effective Prompt Merits via Comparative Evaluation
To systematically analyze effective prompt characteristics, we conduct comparative evaluations. All experiments are implemented using the lightweight LLM $M_{\text{ref}}$ (Qwen2.5-7B-Instruct), which serves as: (1) the merit-guided prompt optimization model, (2) the inference model, and (3) the prompt rewriting model. Below are the details for each evaluation step.
@@
We begin by using a rewriting prompt (shown in Fig. 6) to generate multiple reformulations of each raw question. Specifically, five rewritten versions are produced for each input prompt to preserve diversity. Below is an example from the Alpaca dataset, where the raw prompt results in an incorrect response, while Rewrite 1 and Rewrite 2 yield correct answers under the same inference model:
@@ Appendix C POP Dataset Construction and Training Details
In dataset construction and prompt optimizer training, the lightweight LLM $M_{\text{ref}}$ (Qwen2.5-7B-Instruct) serves as: (1) the prompt degradation model, (2) the inference model, (3) the merit-guided prompt optimization model, and (4) the base model for prompt optimization learning.
@@
To simulate noisy user inputs—such as “how can i go fr sigapor?”—we degrade 10% of the raw prompts from the Alpaca and BPO datasets using a base model. The prompt used for degradation is shown in Fig. 7. Below is an example from the Alpaca dataset:
@@
We fine-tune MePO based on Qwen2.5-7B-Instruct2 using 4 NVIDIA RTX A5000 GPUs. Each input instruction $x$ is truncated or padded to 2000 tokens. Training hyperparameters and cost are summarized in Table 6.
@@
LLaMA3.2-3b-instruct - 73.32 55.46 63.23 35.60 66.32 58.79
Inference Model 78.16 60.75 63.00 37.91 69.37 61.84
BPO 74.75 55.74 56.71 37.34 67.14 58.34
FIPO 78.62 60.98 65.13 37.65 69.53 62.38
MePO 79.55 61.43 65.66 40.49 70.29 63.48
@@
Appendix D Dataset Information and Experimental Implementation
Evaluation Protocol.
GSM8K, PiQA, and BBH are evaluated in a 3-shot setting to facilitate answer extraction, while all other datasets are evaluated in a zero-shot setting. For consistency, all BBH tasks are reformatted into multiple-choice format, except for multistep_arithmetic_two, object_counting, and word_sorting, which retain their original formats. PiQA, which includes two-solution questions, is also converted to multiple-choice format.
Unlike previous Qwen and LLaMA evaluations (Yang et al., 2024; Grattafiori et al., 2024), which select the option (e.g., A, B, or 1, 2) based on the highest next-token logit probability, we extract the answer from the model’s generated reasoning. Specifically, we use format-specific instructions—e.g., for zero-shot ARC: “Reply with the answer option starting with ##, like ##A, ##B, ##C, or ##D”—and extract the answer that follows the ## marker in the response. This approach better reflects the model’s reasoning ability during answer selection.
Prompt Templates.
Fig. 12 illustrates the prompt formats used for GSM8K, PiQA, and ARC. For BBH, multiple-choice tasks follow the PiQA format, with in-context examples prepended to the test query along with the corresponding golden answers. Math-related tasks in BBH adopt the GSM8K format.
Baseline Implementations.
@@ D.1 Case Study: Downward Compatibility
As shown in Table 7, both BPO and FIPO exhibit performance degradation when paired with lightweight inference models. For instance, under LLaMA-3.2-3B-Instruct, compared to using the inference model itself as the optimizer, BPO drops by 3.41% on ARC-Easy, 5.01% on ARC-Challenge, 6.29% on GSM8K, and 2.23% on PiQA. Under LLaMA-3.2-1B-Instruct, FIPO drops by 3.60% on ARC-Easy, and BPO drops by 4.39% on GSM8K. Both also show decreased performance on BBH across the two lightweight inference models. In contrast, MePO consistently improves performance across all datasets, demonstrating strong downward compatibility and robustness in low-resource settings.
@@ D.2 Case Study: Online-Based Discrete APO Comparison
To further validate MePO’s effectiveness in optimizing prompts or instructions with refined meta-prompts, we compare MePO with three discrete APO methods on GSM8K:
@@
Table 8: GSM8K meta-prompts used for each baseline discrete APO method.
@@ D.3 BBH Experimental Results
LLaMA3.3-70b
-instruct
date_understanding 44.8 30.8 28.8 31.6 42 25.6 27.2 62 60.8
disambiguation_qa 63.2 32.8 31.2 32.4 30 36.4 33.6 68.8 65.6
hyperbaton 75.6 48.4 48.4 48.4 48.4 51.6 48.4 78 87.6
logical_deduction_five_objects 49.2 22.4 18 28 33.2 24.4 19.6 50 58.8
logical_deduction_seven_objects 42.4 17.6 14.8 30 27.2 30.4 15.2 42.8 54.4
logical_deduction_three_objects 66.4 32.8 31.6 33.6 36.4 35.2 31.6 71.2 75.6
movie_recommendation 48 26.8 22.4 33.2 23.6 23.2 22.4 52.8 76.4
penguins_in_a_table 54.11 21.23 22.6 31.51 46.58 26.71 23.97 50.68 67.12
reasoning_about_colored_objects 47.6 32.4 32 16.4 29.2 30 16.4 46 51.6
ruin_names 37.2 24 30.8 35.2 27.2 41.6 28.4 66.8 51.2
salient_translation_error_detection 32 26.8 23.2 21.6 28 34 16.8 52 71.6
snarks 60.11 44.94 46.07 46.07 50.56 57.3 46.07 81.46 74.16
word_sorting 88 82.4 80 84 80 52 1.2 100 100
temporal_sequences 61.2 19.2 28.4 28.4 33.2 28 28.4 62 83.6
tracking_shuffled_objects_five_objects 28.4 19.2 20 20.8 30.8 20 20 17.2 40.4
tracking_shuffled_objects_seven_objects 13.2 24 14.4 13.6 22 11.2 12.8 13.6 18.4
tracking_shuffled_objects_three_objects 27.2 29.2 24.4 26.8 49.2 30 30.8 28 31.2
causal_judgement 62.57 49.2 51.87 51.34 49.2 47.59 51.87 53.48 79.68
formal_fallacies 54 52.8 53.2 53.6 52.4 51.6 53.2 52 77.2
navigate 57.6 53.6 42 41.6 41.6 41.2 42 52.8 57.2
web_of_lies 46.4 48.8 48.8 49.2 54 44.4 46.4 46.8 68.4
sports_understanding 70 50 46.4 70.4 54 53.6 43.2 45.6 66.8
boolean_expressions 34.8 56.8 52.4 54.8 51.2 48 54 30.4 62.8
multistep_arithmetic_two 33.6 12.4 9.6 36 14 46 16 16.8 85.6
object_counting 36 33.2 32.8 50.4 33.2 48.8 46.8 49.6 90
@@
LLaMA3.3-70b
-instruct
date_understanding 57.77 36.4 35.6 49.2 52.4 34 31.2 50 64.8
disambiguation_qa 66.4 38.4 37.2 48.8 54.8 44.8 40.8 68.8 71.6
hyperbaton 81.6 54.4 50.4 50.4 54.4 56.8 48.4 62 90.8
logical_deduction_five_objects 58.8 27.2 24.4 47.2 42.4 30 23.6 52 66.8
logical_deduction_seven_objects 49.2 29.6 20.4 39.6 36.8 21.2 18.4 51.6 63.2
logical_deduction_three_objects 82 36.8 51.2 42.8 39.2 40.4 34.8 65.6 80.8
movie_recommendation 63.2 38.4 26.8 38.4 29.2 32 27.2 34 77.6
penguins_in_a_table 56.85 28.77 25.34 38.36 54.79 31.51 25.34 51.37 71.23
reasoning_about_colored_objects 56.8 35.6 38 23.6 41.6 26 18.4 58.8 63.2
ruin_names 38 27.6 41.6 42 28.8 45.6 28.4 73.6 59.2
salient_translation_error_detection 43.2 36.8 38.8 28.8 31.2 39.2 23.6 54.8 72
snarks 67.42 53.37 51.12 49.44 66.85 60.11 46.07 83.15 76.97
word_sorting 95.2 92.4 90 92 86.4 84 3.2 100 100
temporal_sequences 38 37.6 39.2 33.2 42 29.2 28.8 66.8 84.4
tracking_shuffled_objects_five_objects 31.2 39.6 33.2 35.2 32.8 20.4 23.2 21.2 28
tracking_shuffled_objects_seven_objects 15.2 30 26.4 16.4 28.4 18.8 15.2 18 30.4
tracking_shuffled_objects_three_objects 33.6 32.4 33.2 33.2 64.8 31.6 31.6 29.6 31.6
causal_judgement 49.2 56.68 51.87 51.79 56.15 62.57 52.41 59.36 80.21
formal_fallacies 51.2 57.2 54.8 53.2 52.8 53.2 53.2 59.2 83.2
navigate 61.6 58.8 44.4 44.8 48.4 47.2 42 60 42.4
web_of_lies 51.2 55.2 51.6 51.6 59.2 51.2 49.6 49.6 71.6
sports_understanding 77.6 67.2 54.8 77.2 61.2 58 46 49.2 75.6
boolean_expressions 48.4 61.2 54 54.4 54 52.8 55.6 36.4 66.4
multistep_arithmetic_two 45.6 14.4 14.4 41.2 16 57.6 18.4 18 84.4
object_counting 39.6 35.2 60.4 55.2 34 60.8 49.6 53.6 86.8